### v8

@@ 4.2 Model Training and Evaluation Metrics
Validation follows the COCO protocol on the validation split: we report $\mathrm{mAP}_{50}$ (AP at IoU =0.50=0.50) and $\mathrm{mAP}_{50:95}$ (AP averaged over IoU thresholds from 0.50 to 0.95 in steps of 0.05). Inference uses class-agnostic non-maximum suppression with an IoU threshold of $0.7$ and at most 300 detections per image; this IoU is unrelated to the evaluation thresholds.
In the task of UAV-based power transmission line defect detection, the performance of different detection algorithms may vary significantly under diverse operational scenarios. The inference capability of a detection model is typically evaluated using several criteria, including: the computational complexity measured in floating-point operations (FLOPs), and a series of key performance metrics such as precision (P), recall (R), average precision (AP), and mean average precision (mAP). These metrics are formally defined as follows:
$P=\frac{TP}{TP+FP}$
Precision (P) measures the proportion of correctly detected defect instances among all predicted defect instances. A higher precision indicates that the model produces fewer false alarms (false positives), which is critical for avoiding unnecessary inspection or maintenance actions.
$R=\frac{TP}{TP+FN}$
Recall (R) reflects the proportion of correctly detected defect instances among all actual defect instances. A higher recall indicates that the model misses fewer true defects (false negatives), which is essential for ensuring the completeness of defect identification in safety-critical applications.
$AP=\int_{0}^{1}P(R)\,dR$
The average precision (AP) is computed as the area under the precision–recall curve. It captures the trade-off between precision and recall across different confidence thresholds, providing a more comprehensive evaluation of the detection performance for a single defect category.
$mAP=\frac{1}{k}\sum_{i=1}^{k}AP_{i}$
The mean average precision (mAP) represents the mean value of AP over $k$ defect categories. It is widely recognized as the most comprehensive metric for evaluating detection models, as it jointly considers both detection accuracy and robustness across multiple defect types.
In addition to these general measures, the COCO evaluation protocol further defines scale-sensitive metrics, such as: $AP_{s}$ and $AP_{m}$, to provide deeper insights into detection performance across different object sizes. Specifically, $AP_{s}$ evaluates the average precision for small objects (with areas less than 32² pixels), while $AP_{m}$ measures the average precision for medium-sized objects (with areas between 32² and 96² pixels). Since the majority of defects in transmission line inspection datasets are small in scale, $AP_{s}$ serves as a particularly critical indicator, directly reflecting the model’s capability to capture subtle and fine-grained defect patterns that occupy only a limited number of pixels. Therefore, the COCO metrics $AP_{s}$ and $AP_{m}$ are employed in this study to assess detection performance across different object scales.
Here, $TP$ (true positives) denotes the number of correctly detected defect instances, $FP$ (false positives) refers to non-defect regions mistakenly identified as defects, $FN$ (false negatives) indicates defect regions that were not detected, and $k$ denotes the total number of defect categories. Together, these metrics provide a balanced overall performance in terms of accuracy, completeness, and overall reliability, which is important for real-time transmission line inspection tasks.
@@ 4.3 Comparison of heatmaps
To intuitively demonstrate the effectiveness of the proposed modules, we generate heatmaps from the baseline RT-DETR model and the model with the proposed modules (EE-ResNet, SPD Convolution, and CSDMAM) on the same input image. These heatmaps illustrate the attention distribution of the models during the recognition process. As shown in Fig.7, subfigures (a), (c), and (e) represent the attention maps from the baseline model corresponding to the modules before integration, while subfigures (b), (d), and (f) show the attention maps after incorporating the proposed modules. It can be observed that the attention distribution after the inclusion of the modules is more fine-grained, and the object contours are more distinct, indicating that the proposed modules enhance the model’s ability to preserve more detailed information.
@@ 4.5 Comparisons With Previous Methods
In this section, we conduct a rigorous comparative study between the proposed TinyDef-DETR and representative state-of-the-art detectors spanning one-stage (e.g., YOLOv5/8/10/11), two-stage/cascade paradigms, and Transformer-based frameworks (e.g., RT-DETR, Conditional-DETR, DINO). All models are evaluated under a unified protocol—identical data splits, input resolution, and training hyperparameters—to ensure fairness.
Before presenting the experimental data, to visually highlight the differences between TinyDef-DETR and other representative detectors, we conduct a case study on four representative images from the CSG-ADCD dataset. To save space and facilitate readability in a scenario where the majority of objects are small objects, we render the detection bounding boxes with filled colors rather than outlines only. This visualization provides a clearer comparison of recognition results across different models, and the detection results are illustrated in Fig. 8.
@@
Table 3 summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, 10, 11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}_{50}$, and $\text{mAP}_{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs. The best values in each column are highlighted in bold for clarity.
@@
In addition, part of the models are implemented with MMDetection, an open-source object detection toolbox based on PyTorch, which is a core component of the OpenMMLab Chen et al. (2019) project. This ensures standardized training pipelines, reproducibility, and fair comparison across different architectures.
It is also worth noting that DINO(Zhang et al., 2022)—evaluated under identical hardware and software using the official OpenMMLab Chen et al. (2019) recipe dino-4scale_r50_8xb2-12e_coco with its default 12-epoch schedule—performs worse than TinyDef-DETR across all reported metrics, while requiring 79.8 hours of training; by contrast, TinyDef-DETR, trained with our default 100-epoch schedule on the same setup, completes in under 20 hours. Although wall-clock time is environment-dependent, these figures reflect realistic, recommended settings, and TinyDef-DETR thus offers a more favorable accuracy–efficiency trade-off for real-world UAV inspection—amusingly, it seems this “longer training, better performance” effect doesn’t always make sense.
To further highlight performance on small objects and defect-specific categories, Table 4 reports additional evaluation metrics. These include Recall and $\text{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss), which are the two most frequently occurring defect types in transmission line inspection scenarios and are therefore selected as representative categories for detailed evaluation, alongside scale-sensitive measures $\text{AP}_{s}$ and $\text{AP}_{m}$. Such results enable a fine-grained comparison of detection performance.
@@
Model Parameters (M) GFLOPs Precision Recall mAP50 mAP50:95
@@
A comprehensive illustration of the trade-off between detection accuracy and computational efficiency is presented in Fig. 9, where the results reported in Table 3 are visualized. The plot highlights the relative positions of different detectors in terms of $\text{mAP}_{50}$ versus GFLOPs, enabling an intuitive comparison of the accuracy–efficiency balance across lightweight YOLO variants, Transformer-based detectors, and the proposed TinyDef-DETR. Notably, TinyDef-DETR achieves an $\text{mAP}_{50}$ of 0.275 with a computational cost of only 65.3 GFLOPs, striking a favorable balance between accuracy and efficiency compared with both lightweight YOLO models and heavier Transformer-based detectors.
Furthermore, Table 4 provides a more detailed view of performance on small objects and defect-specific categories. It reports Recall and $\text{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss) — the two most frequently occurring defect types in transmission line inspection — alongside scale-sensitive metrics $\text{AP}_{s}$ and $\text{AP}_{m}$. These results offer deeper insights into model behavior and enable fine-grained performance comparisons across representative defect categories.
Finally, a holistic comparison of the selected detectors is presented in the form of a radar chart (Fig. 10). The models included — YOLO v5m, YOLO v8m, YOLO 11m, RT-DETR-R18, RT-DETR-x, and TinyDef-DETR — represent both one-stage and Transformer-based detection paradigms. The evaluation encompasses a comprehensive set of indicators, namely Recall, $\text{mAP}_{50}$, $\text{mAP}_{50:95}$, $\text{AP}s$, $\text{AP}m$, Recall of Polluted Insulator (jyzwh), $\text{mAP}{50}$ of Polluted Insulator, and $\text{mAP}{50}$ of Damaged Pole (dgss). For enhanced visual comparability, each metric is normalized to its maximum value.
4.6 Generalization Experiments for the Proposed Methods
Since small objects dominate our dataset, we further evaluate the proposed method on the VisDrone-DET 2019 benchmark, a widely used UAV dataset for object detection in images. VisDrone-DET 2019Du et al. (2019) comprises 8,599 drone-captured images with >540k annotated bounding boxes across 10 general-purpose categories (e.g., pedestrian, person, car, van, bus, truck, motor, bicycle, awning-tricycle, tricycle), split into 6,471/548/1,580 for train/val/test. Importantly, VisDrone does not include transmission-line defect classes. Therefore, our results on VisDrone (Table 5) assess cross-domain transferability rather than in-domain recognition. Owing to its dense object distribution, the high proportion of small targets, and complex real-world conditions, VisDrone provides a rigorous testbed for robustness and generalization. The strong performance we observe thus demonstrates our framework’s general efficacy for small-object detection in UAV imagery, beyond the specific domain of transmission-line inspection, which is practically valuable.

### v9

@@ 4.2 Model Training and Evaluation Metrics
Evaluation is conducted on the test split following the COCO evaluation protocol, where we report mAP50, the mean Average Precision at an IoU threshold of 0.50, and mAP50:95, the mean Average Precision averaged across IoU thresholds from 0.50 to 0.95 in increments of 0.05. During inference, class-agnostic non-maximum suppression (NMS) with an IoU threshold of 0.7 is applied, retaining at most 300 detections per image. Note that this NMS threshold is independent of the IoU thresholds used for AP and mAP computation.
The COCO protocol additionally provides scale-sensitive metrics that illuminate detector performance across object sizes: $AP_{s}$ measures average precision for small objects (area $32^{2}$ pixels) and $AP_{m}$ measures average precision for medium objects (area between $32^{2}$ and $96^{2}$ pixels). Because most defects in transmission-line inspection datasets are small, $AP_{s}$ is especially informative of the model’s ability to detect fine-grained, low–pixel-count anomalies.
We evaluate inference efficiency and model complexity using standard measures: floating-point operations (FLOPs) quantify computational complexity and model size is reported as the number of parameters (Params, in millions). Detection performance is analyzed using precision ($P$), recall ($R$), average precision (AP), and mean average precision (mAP). These metrics are defined as
$P=\frac{TP}{TP+FP},\qquad R=\frac{TP}{TP+FN},$
where $TP$, FPFP, and $FN$ denote true positives, false positives, and false negatives, respectively. The average precision for a single category is computed as the area under the precision–recall curve,
$AP=\int_{0}^{1}P(R)\,\mathrm{d}R,$
and the mean average precision across $k$ categories is given by
$\mathrm{mAP}=\frac{1}{k}\sum_{i=1}^{k}AP_{i}.$
To provide a more fine-grained assessment relevant to transmission-line inspection, we additionally report scale-sensitive metrics ($AP_{s}$, $AP_{m}$) and category-specific results for the two most prevalent defect types: Polluted Insulator (jyzwh) and Damaged Pole (dgss). For these categories we present both recall and $\mathrm{mAP}_{50}$ to highlight the model’s ability to both locate and precisely localize commonly occurring, safety-critical defects. Together, the described computational and detection metrics provide a balanced appraisal of accuracy, completeness, and operational efficiency for real-time UAV-based transmission-line inspection.
Finally, consistent with the above accuracy–efficiency analyses and evaluation protocol, we emphasize the intended deployment mode: post-mission processing of UAV imagery on a server (rather than on-board inference). Using the same $640\times 640$ input as in our comparisons, TinyDef-DETR sustains 152.19 FPS on a single NVIDIA RTX 3090 (24 GB)—that is, 152.19 images/s or 6.57 ms per image ($1000/152.19\!\approx\!6.57$ ms). This measured throughput is well aligned with the accuracy–complexity trade-offs reported above and is operationally meaningful: a typical batch of 1,000 UAV images can be processed in approximately 6.57 s of pure inference time, readily meeting the requirements of server-side, after-mission defect screening for transmission-line inspection.
4.3 Effects of Kernel Size and Frequency-Domain Features on Detection Performance
We quantitatively assessed the impact of the convolutional kernel size in the Large-Receptive-Field branch of CSDMAM (Section 3.4) by sweeping $k\in\{3,7,15,31,63,127\}$. Table 2 summarizes model complexity (Params, GFLOPs) and detection metrics (Precision, Recall, mAP50, mAP50:95, APs, APm), including the newly added configuration $k=127$.
[…]
Expanding to $k=31$ provides the best overall balance. This setting delivers the highest Recall ($0.263$) and the strongest mAP50 ($0.275$) among all candidates, with clear improvements for small and medium objects ($_{s}=0.106, AP$$_{m}=0.110$) at only a marginal computational increase (GFLOPs $\approx 65.3$). We attribute these gains to a receptive field that is large enough to aggregate global line-structure context while still preserving fine detail, aided by the depthwise/separable parameterization and the frequency-guided FSCA path that stabilizes high-frequency cues without amplifying noise.
Further enlarging the kernel beyond this point does not help. At $k=63$, Precision inches up (0.589) but Recall and small/medium APs drop; at $k=127$, Precision peaks (0.627) yet Recall falls to $0.192$ and overall AP degrades ($_{50}=0.209, mAP$$_{50:95}=0.0933$). We observe two coupled effects: (i) oversmoothing of fine structures caused by excessively wide spatial averaging, which suppresses weak, small targets emphasized by FSCA; and (ii) optimization and efficiency penalties—GFLOPs rise from 65.3 (at $k=31$) to 84.1 (at $k=127$), with diminishing returns and a measurable recall deficit on subtle defects.
Taken together, the ablation supports selecting $k=31$ for the Large-Receptive-Field branch of CSDMAM. This configuration yields the most favorable trade-off among global context capture, preservation of fine-grained defect morphology, and computational efficiency in UAV-based transmission-line inspection. Frequency-domain enhancement (FSCA) remains beneficial across settings, but its effect is maximized when paired with a moderately large, not extreme, spatial kernel—precisely the regime at $k=31$.
Table 2: Performance comparison for different kernel sizes in the Large-Receptive-Field branch.
Kernel Size Params (M) GFLOPs Precision Recall mAP50 mAP50:95 APs APm
3 20.4 64.0 0.406 0.206 0.184 0.0866 0.082 0.056
7 20.4 64.1 0.603 0.210 0.224 0.1032 0.089 0.056
15 20.4 64.3 0.496 0.208 0.213 0.0988 0.089 0.091
31 20.5 65.3 0.534 0.263 0.275 0.1187 0.106 0.110
63 20.8 69.0 0.589 0.220 0.244 0.1108 0.098 0.062
127 22.0 84.1 0.627 0.192 0.209 0.0933 0.083 0.058
To comprehensively assess the impact of the FSCA branch and the integration of frequency-domain information, we evaluate three variants of our framework: TinyDef-DETR (w/o), where the FSCA branch and all frequency-domain operations are removed; TinyDef-DETR (Id), where the FSCA branch is retained but replaced with a pure identity mapping that forwards features without any modulation or frequency-domain interaction; and TinyDef-DETR (Full), which incorporates the complete FSCA-based dual-domain enhancement. The TinyDef-DETR (Id) variant is deliberately constructed as a neutral-control configuration. It is analogous to a biomedical trial investigating the effect of a soft drink on human health, in which one group receives the soft drink while another receives an equivalent volume of plain water. In our setting, TinyDef-DETR (Id) maintains the architectural footprint and computational cost of the FSCA branch while eliminating its adaptive behavior, thereby enabling a clean isolation of the genuine contribution of FSCA-based modulation from the mere presence of an additional branch.
As reported in Table 3, TinyDef-DETR (Full) achieves the most favorable trade-off between detection accuracy and computational complexity. Compared with TinyDef-DETR (w/o), it consistently improves both Precision and Recall, with mAP50 rising from 0.206 to 0.275 and mAP50:95 increasing from 0.0926 to 0.1187, while adding only negligible overhead in parameters and GFLOPs. Although TinyDef-DETR (Id) attains the highest Precision (0.597), its mAP50 (0.232) and mAP50:95 (0.1015) remain clearly below those of TinyDef-DETR (Full). This discrepancy indicates that simply introducing an identity shortcut under the same computational budget cannot explain the observed performance gains; instead, they stem from the adaptive dual-domain feature modulation enabled by the FSCA mechanism.
The per-class and scale-aware results in Table 4 further reinforce this conclusion. For challenging categories such as jyzwh and dgss, TinyDef-DETR (Full) achieves higher Recall and mAP50 than both TinyDef-DETR (w/o) and TinyDef-DETR (Id), indicating enhanced robustness to subtle, low-contrast, and visually ambiguous defects. In addition, TinyDef-DETR (Full) obtains the highest APs and APm, confirming that the proposed dual-domain attention mechanism is effective in preserving fine-grained structural cues for small defects while simultaneously capturing richer contextual information for medium-scale targets. By contrast, TinyDef-DETR (Id) does not exhibit comparable improvements, which substantiates that the performance gains are attributed to the FSCA-based frequency-domain enhancement rather than to increased model capacity or superficial architectural modifications.
Table 3: Comparison of TinyDef-DETR variants with and without FSCA and frequency-domain enhancement. TinyDef-DETR (w/o): w/o FSCA; TinyDef-DETR (Id): Identity mapping.
Model Params (M) GFLOPs Precision Recall mAP50 mAP50:95
TinyDef-DETR (w/o) 20.4 64.3 0.465 0.208 0.206 0.0926
TinyDef-DETR (Full) 20.5 65.3 0.534 0.263 0.275 0.1187
TinyDef-DETR (Id) 20.5 65.3 0.597 0.222 0.232 0.1015
Table 4: Per-class Recall and mAP50 for key defect categories and scale-wise AP. TinyDef-DETR (w/o): w/o FSCA; TinyDef-DETR (Id): Identity mapping.
Model Recall (jyzwh) mAP50 (jyzwh) mAP50 (dgss) APs APm
TinyDef-DETR (w/o) 0.369 0.337 0.389 0.081 0.060
TinyDef-DETR (Full) 0.397 0.391 0.419 0.106 0.110
TinyDef-DETR (Id) 0.377 0.365 0.422 0.089 0.067
@@ 4.4 Comparison of heatmaps
To intuitively demonstrate the effectiveness of the proposed modules, we generate heatmaps from the baseline RT-DETR model and the model with the proposed modules (EE-ResNet, SPD Convolution, and CSDMAM) on the same input image. These heatmaps illustrate the attention distribution of the models during the recognition process. As shown in Fig.7, subfigures (a), (c), and (e) represent the attention maps from the baseline model corresponding to the modules before integration, while subfigures (b), (d), and (f) show the attention maps after incorporating the proposed modules. It can be observed that the attention distribution after the inclusion of the modules is more fine-grained, and the object contours are more distinct, indicating that the proposed modules enhance the model’s ability to preserve more detailed information. For better readability and to highlight the subtle differences, zoom-in views are provided, overlaid on the original images, allowing readers to clearly observe the enhanced attention patterns around key object regions.
@@ 4.6 Comparisons With Previous Methods
In this section, we conduct a rigorous comparative study between the proposed TinyDef-DETR and representative state-of-the-art detectors, including one-stage models (e.g., YOLOv5/8/10/11), two-stage/cascade detectors, and Transformer-based frameworks (e.g., RT-DETR, Conditional-DETR, DINO). All models are evaluated under a unified protocol—identical data splits, input resolutions, and training hyperparameters—to ensure a fair and reliable comparison.
Before presenting the quantitative results, we perform a qualitative case study on four representative images from the CSG-ADCD dataset to visually compare TinyDef-DETR with competing detectors. Since most targets in this dataset are small objects densely distributed in local regions, each image is cropped to focus on the areas with high object concentration, thereby providing a clearer and more informative visualization. To further enhance readability under such crowded conditions, detection results are rendered as filled regions rather than simple bounding box outlines. As illustrated in Fig. 8, TinyDef-DETR yields more accurate and consistent localization of small and densely packed objects compared with DINO, RT-DETR-R18, and YOLOv5s, clearly demonstrating its superiority in challenging small-object detection scenarios.
@@
Table 6 summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, 10, 11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}{50}$, and $\text{mAP}{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs. The best values in each column are highlighted in bold for clarity. Note that “YOLO 11s-P2” denotes a variant in which a P2 feature map was added to the model during training, and a dash (“-”) indicates that the corresponding measurement is not available.
@@
It is worth noting that DINO (Zhang et al., 2022)—evaluated under identical hardware and software configurations using the official implementation with its default 12-epoch schedule—performs worse than TinyDef-DETR across all reported metrics, while requiring 79.8 hours of training. In contrast, TinyDef-DETR, trained for 100 epochs under the same setup (following the canonical schedule adopted in RT-DETRZhao et al. (2024)), completes in under 20 hours. Although wall-clock time is inherently environment-dependent, these numbers correspond to realistic, recommended settings. Importantly, the 12-epoch schedule for DINO and the 100-epoch schedule for TinyDef-DETR (inherited from RT-DETR) are the prescribed defaults; we therefore regard these schedules as intrinsic to each method and adopt them to compare truly “native” models. Under this protocol, TinyDef-DETR delivers a substantially more favorable accuracy–efficiency trade-off for real-world UAV inspection—amusingly, it seems this “longer training, better performance” effect doesn’t always make sense.
@@
Model Params (M) GFLOPs Precision Recall mAP50 mAP50:95
@@
A comprehensive illustration of the trade-off between detection accuracy and computational efficiency is presented in Fig. 9, where the results reported in Table 6 are visualized. The plot highlights the relative positions of different detectors in terms of $\text{mAP}_{50}$ versus GFLOPs, enabling an intuitive comparison of the accuracy–efficiency balance across lightweight YOLO variants, Transformer-based detectors, and the proposed TinyDef-DETR. Notably, TinyDef-DETR achieves an $\text{mAP}_{50}$ of 0.275 with a computational cost of only 65.3 GFLOPs, striking a favorable balance between accuracy and efficiency compared with both lightweight YOLO models and heavier Transformer-based detectors. It is worth mentioning that while YOLO 11s-P2, which adds a P2 feature map to the original YOLO 11s, shows a moderate improvement in precision and recall (e.g., mAP50 increases from 0.155 to 0.181), this simple addition of a P2 layer alone does not lead to a substantial boost in overall detection efficiency or accuracy–efficiency trade-off. This suggests that architectural modifications need to be more carefully designed to achieve meaningful gains, particularly for small-object detection in UAV inspection scenarios.
Furthermore, Table 7 provides a more detailed view of performance on small objects and defect-specific categories. It reports Recall and $\mathrm{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss) — the two most frequently occurring defect types in transmission line inspection — alongside scale-sensitive metrics $\mathrm{AP}_{s}$ and $\mathrm{AP}_{m}$. These results offer deeper insights into model behavior and enable fine-grained performance comparisons across representative defect categories. Although the dataset contains only 13 large objects, which is insufficient for robust evaluation, they are still part of transmission line defects. Our model achieves an $\mathrm{AP}_{l}$ of 0.454 for large objects, representing the average over IoU thresholds from 0.50 to 0.95, demonstrating its capability in detecting large-scale defects despite limited samples.
Finally, a holistic comparison of the selected detectors is presented in the form of a radar chart (Fig. 10). The models included — YOLO v5m, YOLO v8m, YOLO 11m, RT-DETR-R18, RT-DETR-x, and TinyDef-DETR — represent both one-stage and Transformer-based detection paradigms. The evaluation encompasses a comprehensive set of indicators, namely Recall, $\mathrm{mAP}_{50}$ , $\mathrm{mAP}_{50:95}$, $\mathrm{AP}_{s}$, $\mathrm{AP}_{m}$, Recall of Polluted Insulator (jyzwh), mAP50{50} of Polluted Insulator, and $\mathrm{mAP}_{50}$ of Damaged Pole (dgss). For enhanced visual comparability, each metric is normalized to its maximum value.
@@
4.7 Generalization Experiments for the Proposed Methods
To further verify the generalization ability of the proposed method for small-object detection, we conduct experiments on VisDrone-DET 2019, a publicly available and generic UAV object detection benchmark rather than a domain-specific dataset. VisDrone-DET 2019 Du et al. (2019) contains 8,599 drone-captured images with over 540k annotated bounding boxes across ten general-purpose categories (e.g., pedestrian, person, car, van, bus, truck, motor, bicycle, awning-tricycle, tricycle), split into 6,471/548/1,580 for train/val/test. Crucially, VisDrone does not include transmission-line defect classes; thus, our results on VisDrone (Table 8) explicitly measure cross-domain transferability rather than in-domain recognition. Owing to its dense object distributions, the high prevalence of small targets, and complex real-world scenes, VisDrone offers a rigorous testbed for evaluating robustness and generalization.
As shown in Table 8, TinyDef-DETR attains leading overall performance with moderate model complexity. Concretely, it consistently outperforms strong baselines across standard detection metrics and shows clear advantages on small objects, while keeping computation and parameters at a practical level.
These results substantiate that the proposed framework generalizes effectively to small-object detection in UAV imagery beyond transmission-line inspection, indicating practical applicability across domains with different object semantics.